OpenStudy (anonymous):

how would i solve this problem 1/(x+3)+1/x=1/8 i know the final answer is 13+-radical 265 all over2

OpenStudy (campbell_st):

multiply every term by the common denominator 8x(x + 3)| which will give 8x + 8(x+3) = x(x+3) expand and collect like terms 8x + 8x + 24 = x^2 + 3x x^2 - 13x -24 = 0 solve the quadratic use the general quadratic formula \[x = (13 \pm \sqrt{(-13)^2 - 4\times1 \times (-24)})/(2\times1)\] \[x = (13 \pm \sqrt{265})/2\]

OpenStudy (anonymous):

you basically want to flip everything so that you don't have fractions. so what you will do is multiply evertying by all the bases multiply left and right by (x+3) & x & 8

OpenStudy (anonymous):

what he said :)

OpenStudy (anonymous):

don't forget to click on the good answer button the least you can do, since cambell was nice enough to do all of this for you

OpenStudy (campbell_st):

lol.. thanks ks

OpenStudy (anonymous):

:)